# CaPo: Cooperative Plan Optimization for Efficient Embodied Multi-Agent Cooperation

### v1

@@ 5.1 Main results
Performance comparison. We follow CoELA to test two-agent cooperation setting, and compare with classical methods like MHP and RHP, and LLM-driven methods CoELA, ProAgent, and RoCo.
Table 1 summarizes the performance of all compared methods under the two settings of the TDW-MAT task, and shows several observations. 1) Compared with the single-agent baseline RHP, CaPo and all two-agent baselines consistently make significant improvements, showing the effectiveness of multi-agent cooperation in embodied tasks. 2) In multi-agent comparisons, our CaPo outperforms LLM-driven methods by a remarkable margin, e.g., respectively making 16.7% and 8.4% improvement over CoELA and RoCo under the oracle perception setting. 3) CaPo with different LLMs as agent brain exhibits consistent superior performance across all settings. Indeed, CaPo with LLAMA-2 achieves comparable performance with CoELA with GPT-3.5-turbo under oracle perception setting. The improvement of CaPo is derived from its meta-plan and progress-adaptive meta-plan, which both provide strategical and coherent guidance for agent cooperation, thereby improving cooperation performance.
@@
Table 2 reports the performance of all methods on the C-WAH task, and shows similar and consistent observations to those on the TDW-MAT task. Specifically, with GPT-4 agents, our CaPo respectively makes 9.8%, 7.6% and 6.6% relative improvement on CoELA, ProAgent and RoCo, consistently highlighting the superiority of CaPo in enhancing multi-agent cooperation.
@@ 5.2 Ablation Study
Convergence analysis of agent discussion. Here we investigate the convergence of agent discussions, specifically focusing on how many rounds are required for agents to reach a consensus on the meta-plan. In the TDW-MAT environment, we set the maximum number of discussion rounds—referred to as the discussion budget—at three. As shown in Fig. 8, agents reach consensus on the new meta-plan within three rounds in most cases, with 78.9% achieving consensus in the “Stuff” tasks. This meta-plan, which incorporates the states of all agents, enables more efficient cooperation in task completion. Furthermore, by limiting the number of discussion rounds, CaPo strikes a balance between discussion effectiveness and budget, preventing unnecessary or prolonged discussions.
6 Conclusion
In this work, we introduce Cooperative Plan Optimization (CaPo) to enhance cooperation efficiency of LLM-driven embodied agents. CaPo first proposes to create a strategic and coherent meta-plan through multi-turn agents discussion before executing any actions. CaPo first proposes creating a strategic and coherent meta-plan through multi-turn discussions among agents before executing any actions. This meta-plan serves as an action guide to efficiently coordinate multiple agents in completing tasks. During the execution phase, agents dynamically adapt the meta-plan to their latest task progress, maintaining the effectiveness of the meta-plan in coordinating agents to complete tasks efficiently. Experimental results on TDW-MAT and C-WAH tasks show the higher task completion rates and efficiency of CaPo compared to state-of-the-arts.
While CaPo significantly improves multi-agent cooperation efficiency, it has limitations, specifically its heavy reliance on LLMs for reasoning and planning during meta-plan generation and adaptation. As shown in Table 1, agents using stronger LLMs like GPT-3.5 outperform those using weaker ones like LLAMA-2. This dependency is a common challenge for LLM-based frameworks like CoELA.
Reproducibility Statement
We provide detailed descriptions of the two aforementioned embodied environments in Sec.B, covering task settings, as well as the observation and action spaces of the agents. Additionally, we present the detailed prompt designs used in our LLMs in Sec.C of the Appendix. Furthermore, we include a section in Appendix Sec. A.3 to demonstrate the reproducibility of our experimental results on the TDW-MAT environments.
@@ Appendix A Additional Experiments and Discussion
A.1 Baseline models
@@
A.2 Results with GPT-4
Here we further provide results on TDW-MAT task using GPT-4 in Table 5. We can observe that our proposed method CaPo achieves consistently better performance that baseline methods across all tasks, including food and stuff transportation. This also demonstrate the effectiveness of our method on improving multi-agent cooperation.
Method Food ($\uparrow$) Stuff ($\uparrow$) Avg. ($\uparrow$)
RHP 52 49 50
RHP 76 74 75
CoELA 87 83 85
CaPo 90 87 89
Table 5: Transport Rate (TR) comparison on the TDW-MAT task using GPT-4. RHP uses a single agent while all others adopt two agents. Results with oracle perception are reported.

### v2

@@ 5.1 Main results
Performance comparison. We follow CoELA to test two-agent cooperation setting, and compare with classical methods like MHP and RHP, and LLM-driven methods CoELA, ProAgent, and RoCo. Table 1 summarizes the performance of all compared methods under the two settings of the TDW-MAT task, and shows several observations. 1) Compared with the single-agent baseline RHP, CaPo and all two-agent baselines consistently make significant improvements, showing the effectiveness of multi-agent cooperation in embodied tasks. 2) In multi-agent comparisons, our CaPo outperforms LLM-driven methods by a clear margin, e.g., respectively making 4.7% and 5.9% improvement over CoELA and RoCo under the oracle perception setting. 3) CaPo demonstrates consistently superior performance across all settings when paired with different LLMs as the agent brain. Notably, even with a less advanced LLM like GPT-3.5-turbo, CaPo achieves a significantly higher performance compared to the baseline CoELA. For example, under the oracle perception setting, CaPo achieves a transport rate of 84, outperforming CoELA’s 72. The improvement of CaPo is derived from its meta-plan and progress-adaptive meta-plan, which both provide strategical and coherent guidance for agent cooperation, thereby improving cooperation performance.
@@
Table 2 reports the performance of all methods on the C-WAH task. Specifically, with GPT-4 agents, our CaPo respectively makes 9.8%, 7.6% and 6.6% relative improvement on CoELA, ProAgent and RoCo. Interestingly, CaPo, when paired with MHP, where the CaPo agent independently creates and updates the meta-plan, outperforms both a team of two MHP agents and a combination of MHP and CoELA agents. These results consistently highlight the superiority of CaPo in enhancing multi-agent cooperation. We provide more analysis on heterogeneous agents in Appendix A.2.
@@ 5.2 Ablation Study
Convergence analysis of agent discussion. Here we investigate the convergence of agent discussions, focusing on how many rounds are required for agents to reach consensus on the meta-plan. In the TDW-MAT environment, we limit the maximum number of discussion rounds—referred to as the discussion budget—to three (details in Appendix A.6). As shown in Fig. 8, agents reach consensus within three rounds in most cases, with 78.9% achieving consensus in the ”Stuff” tasks. By capping the discussion rounds, CaPo achieves a balance between discussion effectiveness and the budget, avoiding unnecessary or prolonged deliberations. Additionally, we provided an communication analysis in Appendix A.7, which demonstrate CaPo achieves a good balance between cooperation efficiency and communication cost.
6 Conclusion and Limitations
In this work, we introduce Cooperative Plan Optimization (CaPo) to improve the cooperation efficiency of LLM-driven embodied agents. CaPo creates a strategic and coherent meta-plan through multi-turn agent discussions before any actions are executed. This meta-plan acts as an action guide to effectively coordinate multiple agents in task completion. During execution, agents dynamically adapt the meta-plan based on task progress, ensuring its continued effectiveness in coordination. Experimental results on TDW-MAT and C-WAH demonstrate that CaPo achieves higher task completion rates and efficiency compared to state-of-the-art methods.
Potential Limitation: While CaPo significantly improves multi-agent cooperation efficiency, it has certain limitations. Specifically, it relies heavily on LLMs for reasoning and planning during meta-plan generation and adaptation. Additionally, CaPo employs a pre-defined communication protocol for creating and updating the meta-plan, which may restrict its flexibility in dynamic or heterogeneous environments. Addressing these limitations will be a focus of our future work.
Acknowledgments
This work was partially funded by Elekta Oncology Systems AB and a RVO public-private Partnership grant (PPS2102).
@@ Appendix A Additional Experiments and Discussion
A.1 Baseline models and other environments
@@
We conducted experiments on TDW-MAT and C-WAH, which are widely recognized benchmarks for embodied multi-agent cooperation. While other environments, such as ALFRED (Shridhar et al., 2020) and BEHAVIOR-1K (Li et al., 2023a), are notable in this domain, they do not support multi-agent cooperation simulation, and therefore, we did not include them in our experiments. Similarly, works based on Minecraft (Wang et al., 2023b; Cai et al., 2023), such as (Wang et al., 2023c; d), focus on settings that differ from ours, as our framework emphasizes decentralized cooperation with partial observation. Consequently, Minecraft-based environments were also excluded from our experiments.
A.2 Experiments on heterogeneous agents’ cooperation
CaPo introduces a pre-defined communication protocol to create and update the meta-plan, which facilitate the multi-cooperation efficiency. However, it is also very interesting to explore whether CaPo could improve cooperation efficiency when cooperating with other decentralized agents, e.g, RHP and CoELA. Specifically, when cooperating with other agents, the CaPo agent independently creates and update the meta-plan due to the lack of multi-round discussion. We conduct experiments on the TDW-MAT tasks using oracle perception, and the results are shown in Table 5.
Runs Food ($\uparrow$) Stuff ($\uparrow$) Avg. ($\uparrow$)
RHP 0.52 0.49 0.50
RHP+RHP 0.76 (+33%) 0.74(+34%) 0.75(+33%)
RHP+CoELA 0.85 (+40%) 0.77 (+35%) 0.81 (+37%)
RHP+CaPo 0.85 (+38%) 0.80 (+37%) 0.82 (+37%)
CaPo+CoELA 0.86 (+36%) 0.83 (+35%) 0.84 (+35%)
Table 5: Results on heterogeneous agents’ cooperation. Here we report transport rate (efficiency improvement) on TDW-MAT tasks using oracle perception and GPT-4. denotes results from the CoELA paper. The reported EI may fluctuate even for same transport rate, due to slight variations of the performance of baseline agent.
As shown in the above table, when CaPo cooperates with RHP agents, it achieves a higher transport rate compared to both a team of two MHP agents and a combination of MHP and CoELA agents. We attribute this improvement to the independently developed meta-plan, which guides the CaPo agent to complete tasks more efficiently and implicitly enhances multi-agent cooperation efficiency.
Interestingly, when our CaPo cooperates with the more advanced agent CoELA, it further improves the transport rate. This highlights the potential of deploying CaPo to collaborate with other decentralized agents.
@@ A.3 Reproducibility of results
A.4 Results on Additional Environments
To test the generalization and effectiveness of our model on new environments, we add one more experiments on RoCoBench (Mandi et al., 2023), which is a multi-robot collaboration environment. We choose three tasks exhibiting asymmetric observation, which are more relevant to decentralized, partia-observation settings, and the results are shown in the Table 7. Our CaPo framework achieves a higher success rate with fixed steps and requires few steps to complete the task. This further highlights the effectiveness and generalization of our model on embodied multi-agent cooperation.
Sweep Floor Pack Grocery Move Rope
Success rate Step Success rate Step Success rate Step
RoCo $0.95\pm 0.05$ 7.1 $0.44\pm 0.06$ 9.9 $0.65\pm 0.11$ 2.5
CaPo (ours) $0.98\pm 0.13$ 6.5 $0.63\pm 0.11$ 7.6 $0.74\pm 0.17$ 2.2
Table 7: Efficiency comparison on RocoBench (Mandi et al., 2023). We report two metrics, success rate of the task, and average steps required to complete the task.
A.5 Number of discussion round
To generate or update the meta plan, all agents are involved in multi-round discussion. We set the discussion budget to 3 rounds to balance the consensus rate and overall efficiency. As shown in the Table 8, with a 3-round limit, the failure rate decreases to 21.1%, and overall efficiency improves to 84%. Extending the discussion budget beyond 3 rounds offers only marginal benefits in reducing failed cases (e.g., a reduction to 17.8% with 4 rounds) while providing no additional efficiency gains, as efficiency stabilizes at 84%. These results demonstrate that a 3-round discussion budget is a reasonable choice, effectively balancing communication costs and task performance.
Maximal Round 1 2 3 4 5
Rate of failed cases (%) 79.1 53.5 21.1 17.8 15.5
Transport rate (%) 79 82 84 84 83
Table 8: Failed case rate and transport rate under different maximal rounds of discussion.
A.6 Communication cost
For communication cost and efficiency, we want to discuss them from several aspects. Firstly, the task cost mainly includes agent moving cost and communication cost. By comparison, moving cost is often much higher than communication cost, since current robots lack the ability of humans, and their movement is costly especially in complex and large-scale environments like urban areas; while the latter only needs LLM’s inference, and massage sending and receiving, and is indeed much easier and efficient. For example, moving from one room to another can often require much more time than LLM’s inference.
Moreover, to reduce communication cost, our CaPo adopts two strategies. a) agent’s progress-adaptive planning module will evaluate whether meta plan is suitable for current task progress: if yes, then all agents will continue their subtasks, and will not communicate, reducing uncommunication costs; b) with the meta plan as a reference, a few round communication is sufficient to arrive at a new plan, since task progress is often smoothly achieved and thus adaptation the meta plan need slightly small changes. Indeed, we set the maximum round of communication as 3 for each discussion. So these strategies differ from CoELA which needs communication at each iteration, while ours does not.
Method # Rounds of Character Frames Cost # Rounds of Time Cost Time Cost
Meta-plan Cost per from Discussion per Meta-plan per Task
Updates Meta-plan Communication per Meta-plan Update
Food tasks 14 1,757 chars 56 frames 2.4 rounds 20.4 sec 26.2 min
Stuff tasks 17 1,462 chars 51 frames 2.1 rounds 22.7 sec 25.6 min
Table 9: Communication cost analysis on the TDW-MAT dataset with oracle perception.
For average rounds of meta-plan updates and average steps for each meta plan discussion, we report the statistics Table 9. Regarding Food tasks on TDW, agents update the meta-plan for an average of 14 times per task, with each discussion requiring 2.4 steps and a total 1,757 characters for sending (i.e., 4 frames since 1 frame can send 500 characters as in CoELA). While this incurs some time cost, the long-term coherent meta-plan results in a substantial transport rate gain. This demonstrates that creating a coherent meta-plan through communication is significantly more time-efficient than relying solely on embodied actions.
A.7 Collaborating with human
To evaluate human-agent cooperation, we follow the methodology of CoELA and conduct experiments on the C-WAH tasks, where the agent Bob (meta-plan evaluator) is controlled by real human participants. We adopt average steps as an efficiency metric and also ask three participants to rate their teammate agent (i.e., CaPo or CoELA) on a 7-point Likert Scale based on three criteria: communication effectiveness, helpfulness, and trust. The results are detailed in Table 10.
Method Average Steps Comm Effectiveness Helpfulness Trust
CoELA 47 $\pm$ 2 5.8 $\pm$ 0.3 / 7 5.5 $\pm$ 0.4 / 7 6.0 $\pm$ 0.1 / 7
CaPo (ours) 44 $\pm$ 1 6.3 $\pm$ 0.1 / 7 6.1 $\pm$ 0.3 / 7 6.3 $\pm$ 0.2 / 7
Table 10: Comparison of CoELA and CaPo on Human Collaboration Metrics.
Efficiency (Average steps). Humans collaborating with CaPo generally achieve higher efficiency, completing tasks with fewer steps, i.e., 44 steps. This is because both agents and humans can avoid redundant actions through the guidance of the proposed meta-plan.
Human Preferences. (1) Communication effectiveness represents the effectiveness of communication between agent and human participants, e.g., whether the agent understands the message and provides useful information. In general, CaPo demonstrates clearer and more reliable communication than CoELA (6.3 vs. 5.8).This improvement is attributed to CaPo’s multi-round discussion mechanism, which creates a long-term, coherent meta-plan, reducing the need for frequent communication. In contrast, in CoELA, agents and humans have to exchange information more frequently, due to the lack of a coherent plan. Finally, agents didn’t show verbose behavior when making plans (as shown in Figure 3), as we constrained the LLM to generate concise plans and messages in prompting. (2) Helpfulness indicates whether the agent helps human participants complete the task faster. By multi-round discussion and creating coherent meta-plans, CaPo could help both agent and human participants to reduce redundant actions in the embodied environments, e.g., aimless exploration. (3) Trust measures how human participants trust their agent teammate. Humans show a stronger preference for trusting CaPo (6.3 vs. 6.0) during the cooperation. This is because CaPo involves humans in the planning process, seeking their input during multi-round discussions to collaboratively create the meta-plan. In contrast, CoELA agents act more independently and sometimes fail to respond to human messages, which weakens trust.
Through human-agent cooperation experiments, we demonstrate that CaPo enhances communication clarity, actively supports human decision-making, and fosters trust by engaging humans in the planning process. These advantages position CaPo as a more favorable teammate for human collaborators, underscoring its potential for real-world cooperative applications.